### v1

# Probing a minimal dark gauge sector via microlensing of compact dark objects

Abstract
We introduce a minimal Dark Standard Model (DSM) consisting of a single spin-0 particle with dark $U(1)$ gauge symmetry, and completely decoupled from the visible sector. Characterized only by the scalar mass $\mu$ and the dark charge $q$, this framework naturally gives rise to a rich phenomenology, including stable solitonic configurations that behave as dark “mini-MACHOs”. We numerically build and evolve these gauged scalar-field solitons, derive their mass-radius relations, and identify a critical charge beyond which no gravitationally bound configurations exist. By combining these results with microlensing surveys that exclude compact objects heavier than the asteroid-mass scale ($M\lesssim 10^{-11}M_{\odot}$), we obtain the constraint $\mu\gtrsim 10\,\rm eV$ for viable configurations, depending on $q$. Our results represent a step forward in showing that purely gravitational observations can constrain the internal parameters of a dark gauge sector, and provide a framework for exploring broader DSM scenarios through future probes such as gravitational wave detections.
I Introduction
The discovery of dark matter (DM) through its gravitational interaction with visible matter has been confirmed at galactic de Blok et al. (2001); Kleyna et al. (2005); Walker et al. (2006); Battaglia et al. (2008) and cosmological scales Spergel et al. (2003); Hinshaw et al. (2013); Aghanim et al. (2020); Naess et al. (2014); de Haan et al. (2016); Lodha et al. (2025); Abdul Karim et al. (2025). The compelling evidence for the existence of dark matter has driven the development of terrestrial experiments pursuing its direct or indirect detection Gaitskell (2004); Gaskins (2016). Terrestrial experimental searches need to assume a coupling of the DM candidate to visible matter beyond a pure gravitational interaction in order to get a detection signal. While this experimental approach may eventually detect dark matter, current observations imply that there is no empirical evidence that DM couples with standard model (SM) particles Meng et al. (2021); Aalbers et al. (2023, 2024); Aprile et al. (2023); Adams et al. (2025); Chang et al. (2025); Barman et al. (2023); Chan et al. (2019); Song et al. (2024); Albert et al. (2024); Li and Han (2025). In this scenario, an even more elusive possibility arises: that multiple dark matter particles could exist, none of them coupled to those of the SM.
We explore the possibility that dark matter may be composed of distinct particle fields preserving local gauge symmetries, mediated by their own gauge mediators, in analogy with the Standard Model (SM). These fields are minimally coupled to gravity but fully decoupled from SM. We refer to this scenario as the Dark Standard Model, DSM. If such DSM exists, several questions arise: i) what is its particle content?; ii) What gauge symmetries dictate the dynamics of DM interactions?; and, most challenging, iii) how could we experimentally probe the DSM and its associated dark symmetries? As a step forward in addressing these questions, we consider a minimal realization of the DSM consisting of a spin-0 particle with an internal $U(1)$ gauge symmetry. This represents a textbook example of a gauge theory that has the lowest number of free parameters: the scalar mass $\mu$ and the charge $q$.
Despite having only two parameters, this DSM has a rich phenomenology. For instance, when $q=0$ and for masses $\mu$ of the order of $10^{-22}$ eV, the model reduces to the Scalar Field Dark Matter (SFDM) scenario, also known as Fuzzy Dark Matter Model, which successfully reproduces the Cold Dark Matter (CDM) at cosmological scales and, at galactic scales, and may alleviate small-scale (central halo regions) CDM tensions Matos et al. (2024); Hui et al. (2017).
While the SFDM scenario focuses on wave-like dark matter behavior, scalar fields can also behave as particle-like, compact configurations. In this scenario, they can also be used within the CDM model, which considers that dark matter is composed of non-interactive particles Navarro et al. (2010), and these particles can be regarded as scalar-field objects. Moreover, microlensing studies Mao (2012); Petač et al. (2022); Alcock et al. (1996) imply that, if the galactic halos are composed of clusters of massive objects, they cannot have a mass larger than $10^{-11}M_{\odot}$, often referred to as mini-MACHOs. Previous studies with scalar field compact objects (see e.g. Hernandez et al. (2004); Barranco and Bernal (2011); Barranco et al. (2013)) showed that galactic halos could be modeled as collisionless ensembles of scalar field MACHOs with masses below than $10^{-7}M_{\odot}$, which was the upper limit at that time.
Within this point of view, we show that for $\mu\simeq 1$ eV, self-gravitating scalar field configurations have masses around $10^{-11}M_{\odot}$ and can play the role of massive compact halo objects (MACHOs). In such a scenario, the halo dynamics would be indistinguishable from that ina CDM halo, since N-body simulations of CDM are insensitive to particle mass granularity smaller than $10^{5}M_{\odot}$ Navarro et al. (2010). These scalar field mini-MACHOS could have formed in the early Universe via a mechanism analogous to the Affleck-Dine Mechanism known in particle physics Kasuya and Kawasaki (2000).
Static, spherically symmetric, self-gravitating systems minimally coupled to a scalar field with $U(1)$ gauge symmetry, known as charged boson stars, have been described in Jetzer and van der Bij (1989); Jetzer (1989); Pugliese et al. (2013); López and Alcubierre (2023). Recently, the dynamics of these self-gravitating configurations was extensively studied in Jaramillo et al. (2025). In this work, we reinterpret these configurations as dark matter structures within the DSM framework. The resulting solutions are everywhere regular and horizonless, representing solitonic equilibria of the model. In contrast, black hole configurations with nontrivial scalar field profiles exist only for self-interacting scalars under specific resonance conditions Herdeiro and Radu (2020); Hong et al. (2020). Such solutions can be continuously connected to the self-interacting (Q-ball–like) boson star sequence when a small event horizon is introduced. However, full 3+1 numerical evolutions show that they develop a non-axisymmetric instability Nicoules et al. (2025), while their horizonless counterparts remain stable Jaramillo et al. (2025). Therefore, within this theory, the viable static gravitational solitons that could serve as MACHO dark matter candidates are the Reissner–Nordström black holes and the gauged boson stars. Throughout the paper we adopt geometric units ($c=1=G$) except where stated otherwise. In this units, the coupling parameter $q$ then has units of inverse length and can be normalized by the scalar field mass $\mu$, so that ratio $q/\mu$ becomes dimensionless.
II Gauge dark matter solitons
The action for the minimal DSM model is
[…]
where $\mu$ is the mass of the spin-0 particle, $F_{\alpha\nu}=\partial_{\alpha}A_{\nu}-\partial_{\nu}A_{\alpha}$, $g_{\alpha\nu}$ is the spacetime metric, $R$ the Ricci scalar. Here $A_{\nu}$ is the four-potential. Notice that we use the gauge invariant covariant derivative operator defined as $\mathcal{D}_{\alpha}=\nabla_{\alpha}+i\,q\,A_{\alpha}$, where $q$ is the gauge coupling constant or the scalar field charge Jaramillo et al. (2025).
Variation with respect to $g_{\alpha\nu}$ leads to the Einstein field equations, with the stress energy tensor given by
$\begin{split}T_{\alpha\nu}=&2\mathcal{D}_{(\alpha}\Phi\,\mathcal{D}_{\nu)}\Phi^{*}-g_{\alpha\nu}\left(\mathcal{D}^{\beta}\Phi\,\mathcal{D}_{\beta}\Phi^{*}+\mu^{2}|\Phi|^{2}\right)\\ &+F_{\alpha\sigma}F_{\nu\lambda}g^{\sigma\lambda}-\frac{1}{4}g_{\alpha\nu}F_{\sigma\beta}F^{\sigma\beta}.\end{split}$ (2)
Variation with respect to $A_{\alpha}$ leads to the Maxwell equations sourced by the spin-0 current $J^{\alpha}$,
$\nabla_{\nu}F^{\alpha\nu}=J^{\alpha}:=qJ_{\Phi}^{\alpha}\,$, (3a)
$J^{\alpha}_{\Phi}:=i\left({\Phi}^{*}\mathcal{D}^{\alpha}\Phi-\Phi\mathcal{D}^{\alpha}{\Phi}^{*}\right),$ (3b)
here $J_{\Phi}^{\alpha}$ is the conserved current that arises from the U(1) gauge invariance of Eq. (1). Finally, variation with respect $\Phi$ leads to the Klein-Gordon equation
$\mathcal{D}^{\alpha}\mathcal{D}_{\alpha}\Phi=\mu^{2}\Phi\,.$ (4)
The minimal energy, self-gravitating realization of model in Eq. (1) is a static spherically symmetric spacetime with a stationary scalar field spherical solutions, [^1] Often refered as charged boson stars (see e.g.Jetzer and van der Bij (1989)). characterized by a harmonic time dependence of the scalar field of the form $\Phi=\phi(r)e^{i\omega t}$ and an “electric” potential $A_{\alpha}dx^{\alpha}=V(r)dt$. We choose coordinates in which the spacetime metric takes the form
$ds^{2}=-e^{2F_{0}(r)}dt^{2}+e^{2F_{1}(r)}\left(dr^{2}+r^{2}d\Omega^{2}\right)\,$, (5)
with $d\Omega^{2}=d\theta^{2}+r^{2}\sin^{2}\theta\,d\varphi^{2}$ the solid angle. Notice that the metric is isotropic.
Under these assumptions for the spacetime and the spin-0 and spin-1 fields, the system reduces to four coupled ordinary differential equations for the functions $V(r)$, $\phi(r)$, $F_{0}(r)$ and $F_{1}(r)$, see e.g. Appendix B in Jaramillo et al. (2025). We solve the system numerically, noting first that the mass of the scalar field $\mu$ can be factored out as the characteristic length scale of the system, and all dimensional quantities rescale with appropriate powers $\mu$. We solve the system using a spectral method with a compactified radial coordinate $r$ and a Newton-Raphson algorithm, as described in Alcubierre et al. (2022), and employ 24 spectral coefficients for all solutions presented.
Figure 1: Gauge dark matter solitons sequence of solutions. Top panel: Mass vs. radius diagram for families of configurations with different values of $q$. Dashed-line displays the maximum mass configurations. Lower panel: Maximum mass configuration as a function of $q$ in the region where gravitationally bound, and stable configurations exist. Beyond the critical $\tilde{q}_{\rm crit}=1/\sqrt{2}\approx 0.707$, no gravitationally bound solutions are found.
In the process, the eigenfrequency of the the spin-0 field $\omega$ is determined as part of the solution. The equilibrium configurations obtained in this scenario form sequences characterized by the coupling constant $q$, which has an upper bound of approximately $q=\sqrt{4\pi}\mu$ that is limited by Coulomb repulsion Jetzer and van der Bij (1989); Pugliese et al. (2013); López and Alcubierre (2023); Jaramillo et al. (2023); Kleihaus et al. (2009); Kumar et al. (2014). Following previous literatureJetzer and van der Bij (1989); López and Alcubierre (2023), we present results in terms of the dimensionless rescaled charge $\tilde{q}=q/(\sqrt{8\pi}\mu)$, for which the critical value is $\tilde{q}_{\rm crit}=1/\sqrt{2}$. For the solitons considered here, the central amplitude of the scalar field parametrizes each sequence of solutions. The top panel of Fig. 1 shows some sequences for several values of $q$. Here $M$ denotes to the total mass of the configuration, obtained from the asymptotic behavior of the metric coefficients:
Table 1: Gauge dark matter solitons: configurations of maximum mass as a function of the rescaled coupling $\tilde{q}$. We list the dimensionless rescaled charge $\tilde{q}$, the eigenfrequency frequency of the the spin-0 field $\omega$, the maximum value of the scalar field at the origin $\phi(r=0)$, the mass $\mu M$, the total electrical charge $\mu Q_{E}$, and the radius $\mu R_{s}$ of the configurations. The critical value $\tilde{q}_{\rm crit}=1/\sqrt{2}\approx 0.707$ marks the limit beyond which no gravitationally bound solutions exist.
$\tilde{q}$ $\omega/\mu$ $\phi(r=0)$ $\mu M$ $\mu Q_{E}$ $\mu R_{s}$
0 $0.8530$ $0.0765$ 0.6330 0 3.816
$0.7$ $0.9958$ $0.0158$ 5.203 30.96 22.86
$0.705$ $0.9987$ $0.00871$ 9.567 57.69 41.71
$M=-\lim_{r\to\infty}\left(r^{2}\frac{dF_{1}}{dr}\right)\,$, (6)
from the $g_{rr}$ component, or
$M=\lim_{r\to\infty}\left(r^{2}\frac{dF_{0}}{dr}\right)\,$, (7)
from the $g_{tt}$ component. Both should coincide and we use this two quantities to asses accuracy of the code. The first definition corresponding to the Arnowitt–Deser–Misner (ADM) mass, while the second corresponds to the Komar mass (see e.g. Ashtekar and Magnon-Ashtekar (1979); Gourgoulhon (2010); Jaramillo et al. (2023)). Following Alcubierre et al. (2022), the radius of the configuration is defined as the areal radius $R$ where $C:=M(R)/R=e^{-F_{1}}M(r)/r$ is maximized. Here $M(r)$ denotes the Misner-Sharp function, which in isotropic coordinates is $M(r)=-r^{2}e^{F_{1}}\partial_{r}F_{1}\left(1+r\partial_{r}F_{1}/2\right)$. The conserved charge $Q$ associated with the conserved current $J_{\Phi}^{\alpha}$ in Eq. (3b), is given by
Figure 2: Summary of current microlensing limits on MACHO dark matter fraction as a function of lens mass. The solid black lines mark the $95\%$ confidence level exclusion boundaries from recent surveys: Niikura et al. Niikura et al. (2019) covers the sub-lunar window around $10^{-11}-10^{-6}\,M_{\odot}$, Smyth et al. Smyth et al. (2020) covers the Earth–Jupiter range $10^{-6}-10^{-3}\,M_{\odot}$, Wyrzykowski et al. Wyrzykowski et al. (2011) covers the stellar range $\sim 10^{-1}-10\,M_{\odot}$, and Blaineau et al. Blaineau et al. (2022) covers the intermediate range $10-10^{3}\,M_{\odot}$. The red-shaded regions within these curves are excluded.
$Q=4\pi\intop_{0}^{\infty}e^{3F_{1}-F_{0}}r^{2}\phi^{2}\left(Vq+\omega\right)\,dr\,$, (8)
which is related to the total “electrical” charge $Q_{E}$ of the soliton through $Q_{E}=q\,Q$. Far from the source, the electric potential decays as
$V(r)=-\frac{Q_{E}}{4\pi r}+\mathcal{O}\left(\frac{1}{r^{2}}\right)\,$, (9)
which we again use to check accuracy of our numerical solutions. We have checked that the relative differences between the ADM and Komar masses, as well as between the two charge definitions, are below $10^{-4}$.
Note that the total charge $Q$ has no definite sign, as it depends on the product of the gauge coupling $q$, which is fixed in sign, and $\omega$, which can be either positive or negative. The Maxwell Eqs. given by Eq. (3) is invariant under $(\omega,A_{\alpha})\rightarrow(-\omega,-A_{\alpha})$, implying that once the parameter $q$ is fixed, a change in the sign of $\omega$ reverses the total electrical charge, as seen from Eq. (8).
Fig. 1 shows that, for each value of $q$, there is a maximum value of $M$ (dashed-line in the top panel) that increases with $q$. Table 1 lists additional properties of the corresponding dark-matter solitons, including their radius $R_{s}$, and total charge $Q_{E}$. Stability analysis, both perturbative Jetzer (1989) and non-perturbative López and Alcubierre (2023); Jaramillo et al. (2025), have concluded that, configurations with radii larger than that of the maximum mass configurations are stable. Therefore, all configurations on the right side of the maximum (see top panel of Fig. 1) remain stable, while those to the left side are unstable: they either collapse to a black hole, relax into a lower-mass equilibrium state, or disperse. The lower panel of Fig. 1 shows $M_{\rm max}$ as a function of the coupling parameter $q$ in the range $[0,0.705]$. Configurations beyond the critical charge value $\tilde{q}_{\rm crit}=1/\sqrt{2}\approx 0.707$ were found to be gravitationally unbound and therefore unstable. In summary, for each value of $q$, there exist a sequence of solutions. If $q<q_{\rm crit}$, a global maximum-mass configuration separates the stable from the unstable branch. These numerical results constrain the dimensionless quantity $\mu M$ as function of $q$ (see lower panel of Fig. 1).
In the next section, we use microlensing limits on MACHOs to further constrain the allowed values of $M$, and consequently the parameters $\mu$ and $q$.
III Constraints on the free $\mu$ and $q$ from microlensing limits
Figure 3: Constraints in the plane $(\mu,q)$ parameter space derived from the non-observation of microlensing events. The shaded region indicates the values of the scalar-field mass $\mu$ and coupling $q$ consistent with current MACHO limits. Larger masses are excluded by Eq. (11).
Current microlensing observations have imposed tight constraints on the fraction of dark matter that may be due to MACHOs. As summarized in Fig. 2, these limits depend on the mass of the compact objects $\mathcal{M}$. Microlensing observations have effectively exclude MACHOs with $\mathcal{M}>10^{-11}M_{\odot}$ as the dominant dark matter. In the sub-lunar to Earth-mass range ($10^{-11}\lesssim\mathcal{M}/M_{\odot}\lesssim 10^{-6}$), the Subaru Hyper Suprime-Cam (HSC) survey of M31 by Niikura et al. Niikura et al. (2019) detected only one candidate event, placing strong upper limits on the MACHO fraction of dark matter and effectively closing the so-called “lunar-mass window”. At lower masses ($10^{-16}\lesssim\mathcal{M}/M_{\odot}\lesssim 10^{-10}$), Smyth et al. Smyth et al. (2020) reanalyzed the same HSC data, incorporating finite-source-size effects. They found that these effects significantly weaken constraints at the smallest masses, suggesting that asteroid-mass MACHOs could still constitute all of the dark matter. At solar masses and above, OGLE observations toward the Magellanic Clouds, particularly those by Wyrzykowski et al. Wyrzykowski et al. (2011), show that the detected microlensing events are consistent with known stellar populations, thereby ruling out MACHOs with $\mathcal{M}\sim 0.1-10\,M_{\odot}$ as more than a small fraction of the dark halo. For heavier objects in the intermediate range ($\mathcal{M}\sim 10-10^{3}\,M_{\odot}$), Blaineau et al. Blaineau et al. (2022) combined MACHO and EROS datasets, finding that such black hole–like MACHOs can account for at most about $15\%-50\%$ of the dark matter, depending on mass. Overall, microlensing surveys have largely ruled out MACHOs as the dominant form of dark matter, leaving room primarily for mini-MACHOs in the asteroid-mass regime.
We model MACHOs using the gauge dark matter solitons described in the previous section and show how microlensing limits constrain the parameters $\mu$ and $q$. For a fixed $q$, we numerically obtain a maximum value of $\mu M$, a consequence of the self-gravitating nature of the solitons in the minimal gauge dark matter mode in Eq. (1) Kaup (1968).
From the above observational constraints, MACHO masses must satisfy
$\mathcal{M}\lesssim 10^{-11}M_{\odot}=1.48\times 10^{-8}{\rm m}\,.$ (10)
Therefore, we require that the maximum mass of the solitonic configurations in Eq.(1) not exceed this bound, i.e.
$M_{\rm max}<1.48\times 10^{-8}\,{\rm m}\,.$ (11)
Otherwise, stable solutions exceeding this bound would have masses already ruled out by current microlensing limits. This condition imposes a constraint on the scalar-field mass $\mu$:
$\mu>6.76\times 10^{7}x(q)\,{\rm m}^{-1}=13.3x(q){\rm eV}$ (12)
where $x(q)=\mu M_{\rm max}$ is obtained from the numerical sequences shown in the lower panel of Fig. 1, or from some specific configurations in Table 1. The corresponding constraints on $\mu$ as function of $q$ are displayed in the top panel of Fig. 3. In the lower panel, we also show the resulting allowed region with the coupling constant given in SI units, using the conversion
$q^{({\rm SI})}=\frac{\sqrt{8\pi G}\hbar}{c^{2}\sqrt{\mu_{0}}}\,\mu\,\tilde{q}=2.167\times 10^{-46}\,\left(\frac{\mu}{\mathrm{eV}}\right)\tilde{q}\,\mathrm{C}\,$, (13)
where $G$ is Newton’s gravitational constant, $\hbar$ the reduced Planck constant, $c$ the speed of light in vacuum, and $\mu_{0}$ the vacuum permeability.
The above results identify a narrow region in the parameter space where gravitationally bound, stable solitons remain compatible with current microlensing constraints. In particular, microlensing non-detections restrict the dark scalar sector to relatively heavy field masses and small gauge couplings, defining a narrow window of viable mini-MACHO configurations within the DSM framework.
IV Conclusions
The dynamics of visible matter on both galactic and cosmological scales have implications for the existence of dark matter through their gravitational interaction. To date, there is no experimental evidence that dark matter interacts with visible matter through any force other than gravity. However, it is possible that the dark matter sector is not composed of a single particle, but instead forms a self-consistent dark analogue of the Standard Model with its own gauge interactions.
We introduced a minimal Dark Standard Model with a dark $U(1)$ gauge symmetry, giving rise to solitonic “mini-MACHOs.” We have presented this scalar field model not as fuzzy dark matter, but within the CDM framework, where dark matter consists of non-interactive particles. We showed that these particles could be composed of a scalar field, whose rest mass must be greater than $10\,\rm eV$, with more restrictive constraints for a non-zero dark charge.
By combining our derived mass-radius relations for these solitons with microlensing constraints on asteroid-mass objects, we obtain a lower bound of $\gtrsim 10\,\rm eV$ on the scalar mass $\mu$. We derived this bound from the non-observation of microlensing events, combined with the general properties of gauged scalar field stars. Our main results are summarized in Fig. 3, where, for each value of the gauge coupling $q$, we identify a corresponding minimum allowed scalar-field mass. In particular, for $q=0$, the bound is of order $10\,\rm eV$.
From Fig. 3, one may conclude that current microlensing observations impose a lower limit on the scalar-field mass $\mu$, while the gauge coupling $q$ is restricted by the theoretical condition $\tilde{q}<\tilde{q}_{\rm crit}=1/\sqrt{2}$, which ensures the existence of gravitationally bound equilibrium configurations. Only sufficiently heavy fields can form compact objects compatible with current surveys. This shifts the viable parameter space of the minimal DSM to $\mu$-values well above the ultralight fuzzy/axion-like regime.
These configurations are everywhere regular and horizonless, forming dynamically stable solitonic equilibria of the model in Eq. (1). In contrast, as mentioned above, black hole counterparts with nontrivial scalar hair arise only in self-interacting theories under specific resonance conditions and are unstable. Therefore, within the considered framework, the viable static gravitational solitons consistent with current astrophysical constraints are the Reissner-Nordström black holes and the gauged boson stars, the latter standing as natural, stable MACHO dark matter candidates. The limits we have derived on the scalar field mass could also be affected by considering additional terms in the scalar potential, such as a self-interaction.
The dark matter model proposed in this work also opens new directions for exploring its potential observational signatures. Collisions or mergers of these dark asteroids could contribute to a stochastic gravitational wave background with distinctive spectral features, as discussed previously (see e.g. Clesse and García-Bellido (2017); Carr et al. (2016); Giudice et al. (2016); Jaramillo et al. (2022); Lira et al. (2025)). Furthermore, the accretion of such dark compact objects onto black holes could endow Reissner–Nordström and Kerr-Newman geometries with an effective dark charge, providing a novel physical interpretation of the charge parameter in these solutions. This perspective could be applied in other contexts where the conventional view has been used, such as to matter in the vicinity of black holes. This possibility motivates further studies of geodesic motion and observable effects in such spacetimes, following the approaches developed in Herrera-Aguilar and Nucamendi (2015); Wang et al. (2022); Villegas et al. (2023).
Within this scenario, there are multiple other possibilities that can be studied with different gauge mediators and dark matter particles. These bounds are expected to hold in more general dark-gauge sectors with additional fields or interactions. Given the opening of new observational tools and experiments that probe the Universe through gravitational interactions, such as gravitational wave observations, it now seems possible to begin the systematic study of a dark matter standard model along the lines described in this work, potentially connecting gravitational phenomena to the internal structure of the dark sector.
Acknowledgements.
This work has been partially supported by the National Key R&D Program of China under grant No. 2022YFC2204603, by the Generalitat Valenciana (grants CIDEGENT/2021/046 and Prometeo CIPROM/2022/49), and by the Spanish Agencia Estatal de Investigación (grants PID2024-159689NB-C21 funded by MCIN/AEI/10.13039/501100011033, PRE2019-087617, and ERDF A way of making Europe). We acknowledge computational resources and technical sup-925 port of the Spanish Supercomputing Network through the926 use of MareNostrum at the Barcelona Supercomputing Cen-927 ter (AECT-2023-1-0006).